The tennis ball has a mass of 5.7 X 10^-2 kg and has an initial speed of 2.0 m/s. Calculate the speed of the ball when it hits the ground. Ignore air resistance.
1
Expert's answer
2018-12-14T14:31:11-0500

Calculate the final velocity of the ball by keeping the kinetic energy of the ball at initial and final equal by the relation as follows:

1/2 mV_i^2=1/2 mV_f^2

Here, m is the mass of the mass of ball,V_i is the initial velocity of the ball and V_f is the final velocity of ball or the velocity at the time of hitting the ground.

Substitute 5.7×〖10〗^(-2) kg for m , 2 m/s for V_i .

1/2(5.7×〖10〗^(-2) kg )(2 m/s)^2=1/2(5.7×〖10〗^(-2) kg)V_f^2

V_f^2=(2 m/s)^2

V_f=2 m/s